QWE Corporation just paid a dividend of $2 a share. The dividends are expected to grow at 10% a year for the next 3 years, and the 5% per year thereafter. The required rate of return on QWE stock is 12%. What is the current price of QWE?

Question 8 options:

Not enough information.

$22.68

$35.58

$28.47

$34.26
Answer #1

D0 = $2

From then, the dividend grows by 10% for next 3 years

D1 = $2 * 1.10 = $2.20

D2 = $2.20 * 1.10 = $2.42

D3 = $2.42 * 1.10 = $2.662

Terminal value = $2.662 * 1.05 / (0.12 -0.05) = $39.93

Price = Present value of cash flows

Price = [$2.2/(1.12)1] + [$2.42/(1.12)2] + [$2.662/(1.12)3] + [$39.93/(1.12)3]

Price = $34.26